OpenStudy (anonymous):

What are the solutions to the following system of equations? y = x2 + 12x + 30 8x − y = 10

OpenStudy (anonymous):

put y=8x-10 in equation 1and solve for x

OpenStudy (anonymous):

8x-10=x^2+12x+30 subtract 8x from both sides and add 10 x^2+4x+40=0

OpenStudy (anonymous):

ok :)